OpenStudy (anonymous):

Mrs. Perrin, the school music teacher, spent twice as much money on sheet music as she did on CDs. The total amount of money that she spent for sheet music and CDs was one third of her monthly music budget of $189. How much money did Mrs. Perrin spend on sheet music?

OpenStudy (kropot72):

Total amount spent on sheet music and CDs = $189/3 Let x be the amount spent on CDs. Then 2x is the amount spent on sheet music. So we can now form an equation to enable us to find the amount x. \[x+2x=\frac{189}{3}\]

OpenStudy (anonymous):

do i divide 189/3

OpenStudy (kropot72):

Yes. Then add x + 2x.

OpenStudy (anonymous):

63+x+2x

OpenStudy (anonymous):

i still dont get it i am so confuesed

OpenStudy (kropot72):

Not quite. x + 2x = 3x 189/3 = 63 3x = 63.....................(1) Now divide both sides of equation (1) to find the value of x. Then multiply the value of x by 2 to find the amount spent on sheet music.

OpenStudy (anonymous):

but whats the answer

OpenStudy (kropot72):

Its best for you to do the steps that I have posted to find the answer.

OpenStudy (anonymous):

so i diveide 3x to both siade right

OpenStudy (kropot72):

No. You divide each side of the following equation by 3. 3x = 63

OpenStudy (kropot72):

\[\frac{3x}{3}=\frac{63}{3}\]

OpenStudy (anonymous):

x=21

OpenStudy (kropot72):

Correct. So the amount spent on sheet music is \[$21\times 2=$\ you\ can\ calculate\]